Precious Metal Mining has $16 million in sales, its ROE is 18%, and its total assets turnover is 3.2x. Common equity on the firm’s balance sheet is 60% of its total assets. What is its net income? Write out your answer completely. For example, 5 million should be entered as 5,000,000. Round your answer to the nearest cent. Do not round intermediate steps.

$  

Hampton Industries had $75,000 in cash at year-end 2015 and $15,000 in cash at year-end 2016. The firm invested in property, plant, and equipment totaling $170,000. Cash flow from financing activities totaled +$210,000. Round your answers to the nearest dollar, if necessary.

  1. What was the cash flow from operating activities? Enter cash outflows with a minus sign.
    $    
  2. If accruals increased by $35,000, receivables and inventories increased by $200,000, and depreciation and amortization totaled $7,000, what was the firm's net income?
    $    
Answer #1
PRECIOUS METAL MINING:
Per Du Pont identity ROE = (Sales/Total assets)*(Net income/Sales)*(Total assets/Equity)
Substituting available values
0.18 = 3.2*(Net income/Sales)*(1/0.60)
Net income/Sales = 0.18*0.6/3.2 = 0.03375
Net income = Sales*0.03375 = 16000000*0.03375 = $      5,40,000
CHECK:
ROE = 3.2*(540000/16000000)*(1/0.6) = 18.00%
HAMPTON INDUSTRIES:
a) Decrease in cash and cash equivalents during the year = 75000-15000 = $          60,000
Change in cash and cash equivalents during the year = Cash flow from operating activities+Cash flow from investing activities+Cash flow from financing activities
Substituting available values, we have
-60000 = Cash flow from operating activities-170000+210000
Cash flow from operating activities = +170000-210000-60000 = $    -1,00,000
b) -100000 [Cash flow from operating activities] = Net income+35000-200000+7000
Net income = -100000-35000+200000-7000 = $          58,000